If the perimeter of a rectangle is 50 feet, and one of the sides is 10 feet, find the length of the other side.

Answer by checkley77(12844)

2*10+2x=50
20+2x=50
2x=50-20
2x=30
x=30/2
x=15 ans. for the other side.